3 over 4 minus k over 3 = 5 over 12
Answer by nyc_function(2741)

3/4 - k/3 = 5/12

LCD = 12

Multiply every term on both sides by the LCD.

9 - 4k = 5

-4k = 5 - 9

-4k = -4

k = -4/-4

k = 1